Given the following, calculate the estimated cost of ending inventory using the gross profit method.

Gross profit on sales 65 % Net purchases $ 5,000
Beginning inventory $ 30,100 Net sales at retail $ 18,100


Estimated cost of ending inventory            $
Answer #1

Cost of Goods available for sale = 30,100 + 5,000

= 35,100

Cost of Goods Sold = 18,100 - 65%

= 6335

Ending inventory = Cost of Goods available - Cost of Goods Sold

= 35,100 - 6335

= 28,765